How many pounds of 50% hydrogen peroxide is equivalent to 150lbs of 35% hydrogen peroxide?
Answer by LinnW(1048)

Find the number of pounds of hydrogen peroxide in 150 pounds of 35% hydrogen peroxide
0.35 * 150 = 52.5 pounds
Set x = pounds of 50% hydrogen peroxide containing 52.5 pounds of hydrogen peroxide
50% * x = 52.5 pounds
(1/2) * x = 52.5
multiply each side by 2
x = 52.5 * 2
x = 105 pounds